OpenStudy (anonymous):

I have three numbers. The biggest one is twice the middle one, and the biggest one plus the middle one is four times the smallest one. The smallest one plus the middle one is two less than the biggest one. What are the numbers?

OpenStudy (anonymous):

\[A > B > C\]\[A = 2B\]\[A + B = 4C\]\[B + C = A - 2\]

OpenStudy (anonymous):

The algorithm proposed by krebante si right for me : A=16 B=8 and C=6